OpenStudy (anonymous):

25% of 180? and how..

OpenStudy (anonymous):

\[25\% = \frac{25}{100} = .25\] 'of' generally means multiplication. So what is \[.25 \times 180?\]

OpenStudy (anonymous):

25%=1/4 since 25 out of 100 is one out of four so your answer should be 45

OpenStudy (anonymous):

Ok thank you, is there a formula thats works for all of them?

OpenStudy (anonymous):

Yes. \[n\% = \frac{n}{100}\] So: \[n\%\text{ of M } = \frac{n}{100}\times M\]

OpenStudy (anonymous):

25 (180)/100 =45

myininaya (myininaya):

just read what you have of means multiply is means equal

OpenStudy (anonymous):

Its 45 :) polpak

OpenStudy (anonymous):

The best way that works for percentages is knowing the base of 10. For example, 10% of something divides a portion by 10 since 10% is 1 out of 10. 1% of something divides a portion by 100 since 1% is 1 out of 100. Once you know that, you can take the percentage of ANYTHING just based on the previous two statements. Hope this helps!!

OpenStudy (anonymous):

Ok.. so since 1% of 100 is 1 then it can continue with 2,3,4 etc..?

OpenStudy (anonymous):

1/4 of 100 is 25, 1/4 of 80 is 20, 25+20 is 45